OpenStudy (anonymous):

During a substitution method, I worked my math and got -15/7 for my x-value. I checked the back of my book and it gave a whole number for x? Did I do something wrong? Problem: x=-2y x=2-4y (Use Substitution Method)

OpenStudy (skullpatrol):

What did you do?

OpenStudy (anonymous):

Woops the problem was x=3y-4 and 2x-y=7 *Sorry* I plotted in the equation x=3y-4 into 2x-y=7. I got 2(-3y-4)-y=7. I distrubted and then got -6y-8-y=7. Then, I combined like terms getting 7y=15. Using the multiplicative property, I then got y=-15/7 as the book says its -2.

OpenStudy (skullpatrol):

What does the question ask for?

OpenStudy (anonymous):

It tells us to solve the following equations using the substitution method. The back of the book wants the ordered pairs from both equations. Examples: x=y-1, y=4-2x ----> (1,2)

OpenStudy (skullpatrol):

$$\huge \text{I got } 2(-3y-4)-y=7.$$ The minus sign /|\ here infront of 3y is not suppose to be there.

OpenStudy (anonymous):

Wow I was blind from that. Thanks! You really saved my time! c:

OpenStudy (skullpatrol):

Thanks for trying to understand .